OpenStudy (anonymous):

What are the coordinates of the foci of the conic section (y+2)^2/16-(x-3)^2/9=1? I keep getting (3,-7) and (3,3) but the only options are A. (3,-2⨥5) B. (-2,⨥√7,3) C. (-2,⨥5,3) D. (-2,3⨥5)

OpenStudy (phi):

The good news is you have the correct focuses btw, what is that funny sign ⨥ supposed to mean ?

OpenStudy (anonymous):

it is supposed to be the plus/minus sign! sorry lol, so if I have the correct focuses how do i find what the coordinates of the foci are?

OpenStudy (phi):

remember ± is used as short-cut because people don't like to type. in other words, each choice A,B,C or D is really two points both points have the same x value but with two different y values. use + to get the one y value and - to get the other y value

OpenStudy (anonymous):

oh, would that be A then since -2+5 = 3 and -2-5 = -7?

OpenStudy (phi):

*** if I have the correct focuses how do i find what the coordinates of the foci are? *** just to be clear, you have the coordinates of the focuses (foci is an ugly word!) the answer is (3,-7) and (3,3) Yes, choice A is short for those two pair of coordinates.

OpenStudy (anonymous):

thank you so much!! I completely blanked and didnt even think to add and subtract the numbers! thanks again :)

OpenStudy (phi):

that is the trouble with math... people try to say as much as they can in as few symbols as possible (because they don't like to type). The ideas are not that hard, but then people make it look complicated with their "secret" codes.

OpenStudy (anonymous):

yes, that is exactly how I feel!

OpenStudy (anonymous):

Do you think you could help me find the length of the transverse axis for (y+2)^2/16-(x-3)^2/9=1 ?

OpenStudy (phi):

I think it is 2a where a^2 is the number dividing the (y_2)^2 term